OpenStudy (anonymous):

three times the reciprocal of a number equals 9 times the reciprocal of 6. find the number

OpenStudy (anonymous):

3 (1/x) = 9 (1/6), agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 3/x = 9/6

OpenStudy (anonymous):

and 9/6 = 3/2, so 3/x = 3/2. What is x?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Thank you